OpenStudy (anonymous):

The slope of the line below is -1. Write the point-slope equation of the line using the coordinates of the labeled point. (7,-7)

OpenStudy (a_clan):

(y-y1)=m(x-x1)

OpenStudy (anonymous):

slope means gradient grd=-1 (x,y) (7,-7) \[change \in y \]----------\[change \in x\]y--7/x-7 y+7/x-7=-1 cross multiply y+7=-x+7 y=-x-7+7 y=-x

OpenStudy (anonymous):

gimmie a answer!

OpenStudy (anonymous):

y=-x